OpenStudy (anonymous):

The circumference of a circle is 65π. In terms of pi, what is the area of the circle? A) 600 π B) 936.5 π C) 950.25 π D) 1056.25 π

OpenStudy (jdoe0001):

\(\bf circumference=2\pi r\qquad 65\pi =2\pi r\impliedby \textit{what is the "r"adius?} \)

OpenStudy (anonymous):

i dunno

OpenStudy (jdoe0001):

well. solve for "r"

OpenStudy (jdoe0001):

say... if a = bc <--- what would be "c"?

OpenStudy (anonymous):

ummmmmmmm i still dunno

OpenStudy (anonymous):

can you pls help me out my math is due in a few hours

OpenStudy (jdoe0001):

well... you may want to brush up your linear simplification first I'd think